# Reduction-induced the Variation of Partial Von Neumann Entropy

### v1

@@ V Conclusion
In this paper, we propose a novel method called Reduction-induced the Variation of Partial Von Neumann Entropy to quantify QE in bipartite mixed states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to bipartite states. This method exhibits minimal computational complexity. Moreover, it has a wide range of applicability. Its intuitive and clear physical representation, combined with straightforward computation, indicates the ease of exploring its specific potential applications. By considering the degree of bipartite QE, along with the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy proves to be a good measure for assessing bipartite mixed-state QE compared to the existing methods. In future studies, the conceptual framework of the Reduction-Induced Variation of Partial Von Neumann Entropy should be further developed to describe and quantify multipartite QE.

### v2

@@ V Conclusion
In this paper, we propose a method called Reduction-induced Variation of Partial Von Neumann Entropy to quantify QE in bipartite mixed states, which is essentially an extension of Partial Von Neumann Entropy for measuring QE from bipartite pure states to any bipartite states especially for any mixed states. This method exhibits minimal computational complexity, making it highly efficient. Moreover, it has a wide range of applicability. Its intuitive and clear physical representation, combined with simple computation and wide range of applications, indicates the ease of exploring its specific potential applications. By considering the degree of bipartite QE, along with the computational complexity and range of applicability, the Reduction-induced Variation of Partial Von Neumann Entropy proves to be a good measure for assessing bipartite mixed-state QE compared to established methods. In future studies, the conceptual framework of the Reduction-Induced Variation of Partial Von Neumann Entropy should be further developed to describe and quantify multipartite QE.